Operating Systems Concepts

Assume we have a system wherein directory entries can hold up to 16 disk block addresses. For files that are not larger than 16 blocks, the 16 addresses are used as an index table for the file. For files that are larger than 16 blocks, the addresses point to indirect blocks which in turn point to 256 file (data) blocks each. Assume further that a block is1024 bytes. Please calculate how large a file could be (in terms is bytes) in such a system.
Answer #1

Let us consider files larger than 16 blocks then the 16 addresses will point to indirect addresses.

So total addresses pointing to a file block = 16*256 = 24*28 = 212

So the maximum file size = 212*210 bytes = 222 bytes.